An ammonium chloride/ammonia buffer solution is to be prepared with a pH of 10.25. Calculate the [NH3(aq)] / [NH4+(aq)] ratio necessary for this buffer solution. The Ka for ammonia = 1.8x10^-5.
Answer #1

Kb of NH3 = 1.8*10^-5

1st find Ka of NH4+

use:

Ka = Kw/Kb

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Ka = (1.0*10^-14)/Kb

Ka = (1.0*10^-14)/1.8*10^-5

Ka = 5.556*10^-10

pKa = - log (Ka)

= - log(5.556*10^-10)

= 9.255

use:

pH = pKa + log {[conjugate base]/[acid]}

10.25 = 9.255+log {[NH3]/[NH4+]}

log {[NH3]/[NH4+]} = 0.9947

[NH3]/[NH4+] = 9.879

Answer: 9.88